Edward left his estate of $210,000 to his two children and his favorite charity. Each child received 3 times the amount left to his favorite charity. How much did each child and the charity receive?
Answer by checkley77(12844)

3x+3x+x=210,000
7x=210,000
x=210,000/7
x=30,000
3*30,000=$90,000 amount received by each child.
Proof:
90,000+90,000+30,000=210,000
210,000=210,000